OpenStudy (anonymous):

Rewrite the equation in vertex form. x^2+10x+16

OpenStudy (anonymous):

You mean f(x) = a(x - h)^2 + k?

OpenStudy (anonymous):

I could help you if you answered my question.

OpenStudy (anonymous):

Do you need to get it into the form f(x) = a(x - h)^2 + k?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2+10x+16 = 0 x^2 + 10x = -16 x^2 + 10x + 25 = -16 + 25 x^2 + 10x + 25 = 9 (x + 5)^2 = 9 f(x) = (x + 5)^2 - 9